OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution

OpenStudy (anonymous):

will fan and medal!!!!

OpenStudy (texaschic101):

-3(x + 3) = -3(x + 1) - 5 --- distribute through the parenthesis -3x - 9 = -3x - 3 - 5 -- combine like terms -3x - 9 = -3x - 8 -- add 3x to both sides -3x + 3x - 9 = -8 -- add 9 to both sides -3x + 3x = -8 + 9 -- combine terms 0 = 1 (incorrect) NO SOLUTION

OpenStudy (anonymous):

thanks:)

OpenStudy (texaschic101):

sure thing :)